OpenStudy (profanityreally):

Verify: sin^2x/[2-2cosx] = cos^2(x/2)

OpenStudy (welshfella):

use the identity cos x = 1 - 2 sin^2 (x/2) substituting sin^2 x / (2 - 2 cos x) = sin^2x / [2 - 2(1 - 2 sin^2 (x/2)] = sin^2 x / 4 sin^2 (x/2) ...........................(1) Now use the identity sin x = 2 sin (x/2) cos (x/2) giving:- sin^2 x = 4 sin^2 (x /2) cos^2 (x/2) and substituting for sin^x in (1) and you have it. 2

likeabossssssss (likeabossssssss):

welcome to os